OpenStudy (anonymous):

Find the solution of 3x+6y=12 when x =2

jimthompson5910 (jim_thompson5910):

Replace x with 2, then solve for y 3x+6y=12 3*2+6y=12 6+6y=12 6y=12-6 6y=6 y=6/6 y=1 So when x = 2, y = 1

OpenStudy (anonymous):

Could you possibly help me with one more? which of the following points lies on the line 3x+4y=14

jimthompson5910 (jim_thompson5910):

what are your choices

OpenStudy (anonymous):

They are -2,5 2,-5 1,-4 -1,4

jimthompson5910 (jim_thompson5910):

each ordered pair is of the form (x,y) so for instance, the point (1,-4) means x = 1 and y = -4

jimthompson5910 (jim_thompson5910):

let's see if that point works or not 3x+4y=14 3*1+4(-4)=14 ... plug in x = 1 and y = -4 3 - 16 = 14 -13 = 14 that's false, so (1,-4) is NOT a solution to 3x+4y=14

OpenStudy (anonymous):

okay, would -2,5 work?

jimthompson5910 (jim_thompson5910):

plug it in and find out, use the steps I showed above

jimthompson5910 (jim_thompson5910):

then tell me what you get